OpenStudy (anonymous):

When adding square root 16 and -8,which type of number is the sum

OpenStudy (alexandervonhumboldt2):

4-8=-4 intenfer

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

integer right

OpenStudy (alexandervonhumboldt2):

y es sqrt16 -8=4-8=-4